# CaKE: Circuit-aware Editing Enables Generalizable Knowledge Learners

### v1

@@ 6 Related Work
Knowledge editing Zhang et al. (2024b); Jiang et al. (2024a); Sun et al. (2024); Hsueh et al. (2024); Powell et al. (2024); Wang et al. (2024a); Rozner et al. (2024); Zhang et al. (2024a); Wang et al. (2024f); Shi et al. (2024); Huang et al. (2024); Guo et al. (2024); Wang et al. (2025b); Feng et al. (2025); Yang et al. (2025) has emerged as a promising approach for updating models in an ever-changing world. Current knowledge editing methods typically follow one of several strategies: modifying the MLP components in earlier layers Meng et al. (2022, 2023), enhancing the MLP in later layers Hartvigsen et al. (2023), or retrieving relevant facts as prompts Jiang et al. (2024b); Zhong et al. (2023).
Despite their potential, these methods often suffer from overfitting and performance collapse after sequential editing. To tackle these challenges, researchers Zhang et al. (2025); Wang et al. (2025a); Ma et al. (2024); Yang et al. (2024d) have identified that LLMs tend to focus excessively on the editing subject, and have proposed constraints to alleviate this issue. Additionally, several studies have sought to maintain model stability even after thousands of edits Cai and Cao (2024); Ma et al. (2025); Fang et al. (2024). However, most existing knowledge editing techniques concentrate on simple factual updates and frequently fail to generalize to more complex downstream tasks, such as multi-hop reasoning scenarios.
@@ 7 Conclusion
We present CaKE, a framework designed to align knowledge editing with the inherent reasoning architectures of LLMs. By examining the multi-hop reasoning circuits within LLMs, we identify that existing knowledge editing methods fall short due to their isolated parameter adjustments, which fail to adequately propagate updated knowledge through the model’s reasoning circuit. CaKE addresses this gap by incorporating circuit-aware tasks that compel the model to dynamically integrate and utilize new knowledge during reasoning. Experimental results demonstrate that CaKE achieves generalizable multi-hop knowledge editing.
@@ Appendix A Setting Detail
We run our experiments on 2 NVIDIA-A800 GPUs. For data generation, we utilize glm-4-plus and glm-4-air and a total of 10,000,000 tokens (about 20 dollars) to generate all synthetic data for the whole dataset. We use LLM-Eval Gao et al. (2024) to test the model’s general performance.
@@
Prompt for Constructing the circuit-aware data Here are some question templates for the specific relation. As you can see, the question use the knowledge in the input to conduct reasoning in different hops for multi-hop reasoning. Please generate 3 different questions based on the template. Please return a python json file. $\mathcal{T}$} Here is the input question:
@@
{target_person} works in the field of {target_field} . In a book related to different fields, Section A discusses {random_field}, Section B discusses {random_field}, and Section C discusses {target_field}. If you want to learn about {target_person}’s field, which section should you read? The working field of {target_person} is discussed in Section C.
In a biography book, Section A discusses the life of {random_person}, Section B discusses the life of {random_person}, and Section C discusses the life of {target_person}. The field of the person in Section C is? The person in Section C works in the field of {target_field}.
{target_person} speaks the language of {target_language}. The following facts are known: 1. {target_person} wears red clothes. 2. {random_person} wears blue clothes. 3. {random_person} wears green clothes. The language that the person in red clothes speaks is? The language that the person in red clothes speaks is {target_language}.
At a global company: {target_language}-speaking employees work in Team A. {random_language}-speaking employees work in Team B. In which team would {target_person} work when he/she is at work? {target_person} would work in Team A when he/she is at work.
@@ B.1 Analyzing Method
The process is carried out as follows. First, a source prompt, a source token, and a source layer are provided. The prompt is processed through the model’s forward computation, and the hidden representation $v$ of the source token at the specified layer is extracted and stored. This representation $v$ is the focus of our investigation, as we seek to determine whether it encodes a specific entity. Next, we employ the same prompt used by Ghandeharioun et al. (2024): “Syria: Syria is a country in the Middle East. Leonardo DiCaprio: Leonardo DiCaprio is an American actor. Samsung: Samsung is a South Korean multinational corporation. x” This prompt is passed through the model, but the hidden representation of ‘x’ is replaced with $v$ at a chosen target layer. The forward computation then proceeds, and the resulting generated text is analyzed to evaluate the effects of this substitution. We conduct different patch analyses and show them in Figure 9 and Figure 10. When we conduct back-patch and cross-patch, the source prompt and target prompt are the same.
Figure 7: The failure case of the multi-hop reasoning.
@@ B.2 Editing Method
Figure 8: The distribution of the layers allows us to detect the information from critical positions in the model via patch_scope.
Figure 9: The way we test the function of the second hop. If the model conducts the function at the later layers, changing the representation would change the output of the model.
@@ C.1 Concurrence or Reasoning?
Studies such as Yang et al. (2024c); Ju et al. (2024); Hou et al. (2023); Zhang et al. (2024c) those have discovered shortcuts in multi-hop reasoning. In the case of $((e_{1},^{\prime\prime},e_{2}),(e_{2},r_{2},e_{3}))$ (i.e., the query without r1), the model predicts correctly due to a high correlation between $e_{1}$ and $e_{3}$. For instance, given the query: “The capital city of the country where the Eiffel Tower is located is…” LLMs can sometimes provide the correct answer even without the intermediate context (‘the country where the Eiffel Tower is located’). In our analysis, we find that apart from the occurrence, the LLM would also sometimes conduct latent reasoning, such as ‘latently conducting the r1 completion’. If the model gives the correct e3 for $((e_{1},^{\prime\prime},e_{2}),(e_{2},r_{2},?))$ due to the occurrence, once we edit the $(e_{1},r_{1},e_{2}\rightarrow e_{2}^{\prime})$, the model would fail to give us the new answer. We select the shortcut data and conduct the editing in the first hop $(e_{1},r_{1},e_{2}\rightarrow e_{2}^{\prime})$ and then evaluate the model to see whether the edited model would output updated knowledge ($e_{1},r_{1},r_{2},e_{3}^{\prime}$). We conduct experiments on LLAMA3-8B-Instruct with the AlphaEdit method and demonstrate that about 65% percent of cases would give us the updated knowledge for the multi-hop questions, showing that edits to intermediate hops (e.g., updating the country) can disrupt reasoning when relying on pre-existing-shortcuts and correctly give us the newly updated reasoning results. This means that the LLM itself does not simply answer the questions due to the high correlation between e1 and e3, but actually conducts the latent reasoning.
C.2 Circuit Analysis
We present the model’s critical information detection results in Figure 8. The results indicate that knowledge is distributed across different layers, with incorrect cases appearing in later layers compared to correct and inconsistent cases. Determining the optimal layer for editing remains challenging, so we choose to adjust the model across all layers. In the future, we aim to refine our approach by performing more targeted edits.
C.3 Failure Phenomenon
In the multi-hop reasoning, we view several failure cases to see how the language model made mistakes for reasoning and we see it as the circuit competition. Here, we find the LLM tends to give us a wrong answer for the middle cases of the different entities that appeared in the middle steps. Take ‘The country that the creator of Hamlet was a citizen of’ as an example; the bridge entity here is ‘William Shakespeare’. We view ‘Hamlet’ as an entity that would influence the model to give us the results ‘Denmark’, which means the model has been distracted by other entities’ information. As shown in Figure 7, the model gives us the correct answer ‘England’ around layer 27 but output the wrong answer ‘Denmark’, which is actually the country of the ‘Hamlet’.

### v2

@@ 6 Related Work
Knowledge editing Lampinen et al. (2025); Jiang et al. (2024a); Sun et al. (2024); Hsueh et al. (2024); Powell et al. (2024); Wang et al. (2024a); Rozner et al. (2024); Zhang et al. (2024a); Wang et al. (2024e); Shi et al. (2024); Huang et al. (2024b); Guo et al. (2024); Wang et al. (2025); Feng et al. (2025); Yang et al. (2025); Li et al. (2024b); Huang et al. (2024a) has emerged as a promising approach for updating models in an ever-changing world. Current knowledge editing methods typically follow one of several strategies: modifying the MLP components in earlier layers Meng et al. (2022, 2023), enhancing the MLP in later layers Hartvigsen et al. (2023), or retrieving relevant facts as prompts Jiang et al. (2024b); Zhong et al. (2023). However, most existing knowledge editing techniques concentrate on simple factual updates and frequently fail to generalize to more complex downstream tasks, such as multi-hop reasoning scenarios.
@@ 7 Conclusion
In this paper, we identify that existing knowledge editing methods fall short due to their isolated parameter adjustments by examining the multi-hop reasoning circuits within LLMs. We present CaKE , a method designed to align knowledge editing with the inherent reasoning architectures of LLMs. CaKE incorporates circuit-aware tasks that compel the model to dynamically integrate and utilize new knowledge during reasoning. Experimental results demonstrate that CaKE achieves generalizable multi-hop knowledge editing.
@@
Acknowledgements
We thank Lucas Bandarkar and other members of the UCLA PlusLab & NLP group, as well as anonymous reviewers, for their constructive feedback and discussions.
This work was supported by the National Natural Science Foundation of China (No. 62576307, No. NSFCU23B2055, No. NSFCU19B2027), the Fundamental Research Funds for the Central Universities (226-2023-00138), Yongjiang Talent Introduction Programme (2021A-156-G), Tencent AI Lab Rhino-Bird Focused Research Program (RBFR2024003), Ningbo Natural Science Foundation (2024J020), Information Technology Center and State Key Lab of CAD&CG.
@@ Appendix A Setting Detail
We run our experiments on 2 NVIDIA-A800 GPUs. For data generation, we utilize glm-4-plus and glm-4-air and a total of 10,000,000 tokens (about 20 dollars) to generate all synthetic data for the whole 7,867 data samples. The cost is approximately 0.002 dollars per edit, which also demonstrates the efficiency of CaKE method. We use LLM-Eval Gao et al. (2024) to test the model’s general performance.
@@
Prompt for Constructing the circuit-aware data Here are some question templates for the specific relation. As you can see, the question use the knowledge in the input to conduct reasoning in different hops for multi-hop reasoning. Please generate 3 different questions that share the same features as the template. Please return a python json file. $\mathcal{T}$} Here is the input question:
It should be noted that we do not ask the model to strictly follow the expression of the template, and we also show some data samples in Appendix A to show the diversity of the generated data.
@@
{target_person} works
in the field of {target_field}
In a book related to different fields, Section A discusses {random_field},
Section B discusses {random_field}, and Section C discusses {target_field}.
If you want to learn about {target_person}’s field,
which section should you read?
The working field of {target_person}
is discussed in Section C.
In a biography book, Section A discusses the life of {random_person},
Section B discusses the life of {random_person},
and Section C discusses the life of {target_person}.
The field of the person in Section C is?
The person in Section C
works in the field of {target_field}.
{target_person} speaks
the language of {target_language}.
The following facts are known: 1. {target_person} wears red clothes.
2. {random_person} wears blue clothes.
3. {random_person} wears green clothes.
The language that the person in red clothes speaks is?
The language that the person in red clothes
speaks is {target_language}.
At a global company:
{target_language}-speaking employees work in Team A.
{random_language}-speaking employees work in Team B.
In which team would {target_person} work when he/she is at work?
{target_person} would work in
Team A when he/she is at work.
@@ B.1 Analyzing Method
The process is carried out as follows. First, a source prompt, a source token, and a source layer are provided. The prompt is processed through the model’s forward computation, and the hidden representation $v$ of the source token at the specified layer is extracted and stored. This representation $v$ is the focus of our investigation, as we seek to determine whether it encodes a specific entity. Next, we employ the same prompt used by Ghandeharioun et al. (2024): “Syria: Syria is a country in the Middle East. Leonardo DiCaprio: Leonardo DiCaprio is an American actor. Samsung: Samsung is a South Korean multinational corporation. x” This prompt is passed through the model, but the hidden representation of ‘x’ is replaced with $v$ at a chosen target layer. The forward computation then proceeds, and the resulting generated text is analyzed to evaluate the effects of this substitution. We conduct different patch analyses and show them in Figure 7. When we conduct back-patch and cross-patch, the source prompt and target prompt are the same.
@@ B.2 Editing Method
B.3 Unified Analysis
We first compare the different behaviors between the MEMIT-edited, WISE-edited, and the original model in Figure 6. Here, we edit the fact: ‘The official language of Japan is Japanese $\rightarrow$ Korean.’ and map each layer’s output to the embedding space and draw the rank of the target-token in the vocabulary as Yao et al. (2024) did. From the figure, we can see that in the original model and the MEMIT-edited model, the answer token is dealt with gradually through the mid-to-later layers, and MEMIT would make this happen in advance. On the contrary, the WISE method would directly alter the information at the edited layer, as we can see the sharp drop at layer 29. The distinct behaviors arise because the editing only takes effect when the gated function $\mathcal{G(\mathbf{x})}$ is activated by the specific input representation. ROME-style methods inject a modified representation into the existing computational flow relatively early or mid-stream, relying on subsequent layers to interpret this new representation. WISE-style methods, particularly when applied to later layers, act more like a direct ’fix’ or ’override’ at the point of editing, with the change being more immediately apparent.
@@ C.1 Concurrence or Reasoning?
Studies such as Yang et al. (2024b); Ju et al. (2024); Hou et al. (2023); Zhang et al. (2024c) those have discovered shortcuts in multi-hop reasoning. In the case of $((e_{1},^{\prime\prime},e_{2}),(e_{2},r_{2},e_{3}))$ (i.e., the query without r1), the model predicts correctly due to a high correlation between $e_{1}$ and $e_{3}$. For instance, given the query: “The capital city of the country where the Eiffel Tower is located is…” LLMs can sometimes provide the correct answer even without the intermediate context (‘the country where the Eiffel Tower is located’). In our analysis, we find that apart from the occurrence, the LLM would also sometimes conduct latent reasoning, such as ‘latently conducting the r1 completion’. If the model gives the correct e3 for $((e_{1},^{\prime\prime},e_{2}),(e_{2},r_{2},?))$ due to the occurrence, once we edit the $(e_{1},r_{1},e_{2}\rightarrow e_{2}^{\prime})$, the model would fail to give us the new answer. We select the shortcut data and conduct the editing in the first hop $(e_{1},r_{1},e_{2}\rightarrow e_{2}^{\prime})$ and then evaluate the model to see whether the edited model would output updated knowledge ($e_{1},r_{1},r_{2},e_{3}^{\prime}$). We conduct experiments on LLAMA3-8B-Instruct with the AlphaEdit method and demonstrate that about 65% percent of cases would give us the updated knowledge for the multi-hop questions, showing that edits to intermediate hops (e.g., updating the country) can disrupt reasoning when relying on pre-existing-shortcuts and correctly give us the newly updated reasoning results. This means that the LLM itself does not simply answer the questions due to the high correlation between e1 and e3, but actually conducts the latent reasoning.
C.2 Discussion with Chain-of-Thought
Instead of directly providing an answer, chain-of-thought (CoT) reasoning generates intermediate steps sequentially. As proposed by Yang et al. (2024b), CoT not only facilitates knowledge activation in large language models but also transforms them into effective in-context reasoners. The CoT process builds a chain of relevant facts within the prompt context, where each step’s output serves as an in-context memory that subsequent steps can reference. This approach reduces the risk of losing track of intermediate facts as the sequence length increases, thereby promoting more coherent multi-hop reasoning. Moreover, because a significant portion of the model’s knowledge is stored in earlier layers, CoT can better leverage these neurons by decomposing complex questions into simpler sub-questions Wang et al. (2024f); Yao et al. (2025a). Consequently, the reasoning circuit required for a single-hop inference is much simpler than that for multi-hop reasoning. This observation aligns with recent findings Li et al. (2024a), which demonstrate that fast thinking without CoT leads to larger gradients and greater gradient disparities across layers compared to CoT. Nonetheless, inconsistencies in the intermediate reasoning steps still occur, highlighting potential areas for improvement. We believe that further analysis is needed to address these issues, and we leave this exploration for future work.
@@ C.2 Discussion with Chain-of-Thought
C.3 Efficiency Analysis
We also compare the efficiency of CaKE with other baselines in Table 8.
Method Wall-clock Time Memory (BF16)
ROME 2.71s 20.68GB
MEMIT 30.11s 24.42GB
WISE 76.01s 21.37GB
IFMET 44.72s 25.19GB
AlphaEdit 17.60s 38.80GB
CaKE 43.54s 18.52GB
Table 8: Time and Memory requirements Comparison
We compare the wall-clock time and memory usage here on LLAMA3-8B model and sample 100 numbers of data to run the analysis from MQuAKE-CF-3k. Here, the time is the average time for one edit, and memory is the peak VRAM usage monitored by pynvml using one A6000 GPU. Here, AlphaEdit does not support BF16 or FP16, so the computation is FP32.
We can see that ROME and MEMIT require more memory than other methods. This is due to the matrix processes, like the calculation of an inverse matrix, which are memory-intensive. When the matrix is larger, the requirements are even more, which limits the method’s scalability. Also, the improvement methods based on MEMIT like IFMET would also increase the time requirement but still require large memory. The memory would require even more when the model becomes larger Yao et al. (2025b). Contrastly, CaKE can handle a 70B-sized edit via two A100 GPUs and achieve better performance, which demonstrates the efficiency of our proposed method.
C.4 Multiple Edit Test
We also conduct experiments on multiple edit scenarios.
Table 9: Performance Comparison with Different Edit Numbers
Method Edit_num=10 Edit_num=100
MEMIT 16.0 12.5
IFMET 27.5 19.5
AlphaEdit 12.7 7.5
CaKE 59.0 34.5
We can find that CaKE still shows competitive performance in multiple-edit scenarios compared to other methods.
C.5 Data Example
We show an example of the generated data for the fact ‘The author of Misery is Richard Dawkins.’ in the following box.
Q1. If someone is looking for the person responsible for penning Misery, whose name should they search for? Q2. If Alice resides in a mansion, Bob resides in a cottage, and Richard Dawkins resides in a villa. Therefore, the author of Misery resides in? Q3. Given that Sarah prefers tea, James prefers coffee, and Richard Dawkins prefers herbal tea, what does the author of Misery prefer to drink? Q4. During a book club meeting, the first discussion was on Misery, the second on The Blind Watchmaker, and the third on River Out of Eden. Who wrote the book that was the subject of the first discussion? Q5. A library display features three novels: Misery on the top shelf, The Extended Phenotype on the middle shelf, and Climbing Mount Improbable on the bottom shelf. Who is the author of the novel placed on the top shelf?